# Large Language Models Hallucination: A Comprehensive Survey

### 09 Oct 2025 v2

@@ 1 Introduction
Furthermore, we expanded the existing taxonomy of hallucination mitigation techniques derived from prior research [8, 9, 10] by categorizing them into four categories: prompt, retrieval, reasoning, and model-centric training and adaptation-based approaches. Prompt-based mitigation approaches depend on structured prompting strategies to guide models towards generating factual content. Retrieval-based mitigation methods depend on external knowledge to ground outputs. Reasoning-based mitigation techniques, such as chain-of-thought prompting (CoT) and self-consistency, enhance logical coherence and internal consistency. Model-centric training and adaptation-based approaches involve modifying architectures, adjusting training objectives, and employing fine-tuning procedures to improve models’ intrinsic factuality and reliability. Based on our analysis, we show that no single approach completely mitigates hallucination. Consequently, more effective mitigation is a combination of complementary techniques. The most promising are hybrid approaches that combine prompting or reasoning-based techniques with retrieval-based and model-centric training and adaptation strategies.
@@
  • Hallucination mitigation discussion: The hallucination mitigation methods have classified in this survey into four main categories: prompt, retrieval, reasoning, and model-centric training and adaptation-based techniques. Each category have been discussed deeply to show its potential for hallucination mitigation.
@@ 2 Related Surveys
Given the rapid evolution of LLMs and their expanding applications across diverse domains, hallucination detection and mitigation techniques have also progressed. Consequently, there is a need to thoroughly analyze the hallucination causes, detection, and mitigation techniques to remain up-to-date. Investigating these factors will contribute to diagnosing the root causes of LLM hallucinations and developing more effective detection and mitigation techniques. Several surveys have been published recently on LLMs hallucination [5, 9, 8, 11, 12]. These surveys examine LLMs’ hallucination from multiple points of view and provide significant insights. Ji et al. [5] reviewed hallucination across various NLG tasks and outlined the mitigation methods applied and the evaluation metrics used in each task. Ye et al. [12] extended this work by proposing a new taxonomy for the detection and mitigation of hallucination. Zhang et al. [11] highlighted some issues related to LLMs, such as input, context, and fact-conflicting hallucinations. Tonmoy et al. [8] focused on hallucination mitigation by presenting a taxonomy that categorizes mitigation techniques into prompt engineering, retrieval-augmented generation (RAG), self-refinement, and decoding strategies. Huang et al. [9] presented a dual taxonomy of factuality and faithfulness hallucinations and defined hallucination causes in the data, training, and inference stages. The survey also linked detection and mitigation strategies to their foundational causes to guide robust system development.
@@ 3.1 Definition of Hallucination
In a psychological context, the term ”hallucination” refers to a false perception of objects or events [19]. It involves the experience of seeing, hearing, feeling, smelling, or tasting stimuli that are not actually present. These sensations are often indistinguishable from reality to the one experiencing them [20]. In the context of LLMs, hallucination refers to the generation of text that appears reasonable, fluent, and coherent but lacks grounding in factual or accurate information [21]. This phenomenon poses a considerable challenge, particularly in domains where precise information is essential, such as healthcare, legal counsel, finance, and education. For instance, in a scientific article summary, a model might introduce non-existent research findings. In healthcare, a hallucinated medical fact could be hazardous since incorrect medical information can mislead patients or healthcare professionals, leading to inaccurate diagnoses, inappropriate treatments, or neglect of critical conditions. Consequently, hallucination compromises the reliability and trustworthiness of LLMs, which poses potential risks in real-world applications.
@@ 3.3 Hallucination in NLG Tasks
Machine Translation involves translating a text from one language to another. Models may provide grammatically accurate translations but may convey new or irrelevant meanings, particularly for low-resource languages or ambiguous inputs [24]. Text Summarization is generating a concise and coherent summary that preserves the most important information from one or more source documents. In the literature, two main methods for text summarization are followed: extractive and abstractive [25]. Extractive text summarization selects and organizes the most important sentences from the source text to form a summary. Abstractive text summarization, on the other hand, paraphrases the selected sentences to form a summary. Abstractive text summarization is more prone to hallucination than extractive text summarization, since paraphrasing the text may lead to generating false information [21].
@@
Generative QA is a type of QA system that generates answers to user queries in natural language, rather than selecting a predefined answer from a database or document. It constructs its response based on the model’s comprehension of the inquiry and the provided context. Generative QA systems search for external knowledge from multiple sources to formulate a response. These sources may consist of redundant, complementary, or conflicting information [26, 27]. Therefore, it is highly susceptible to generating hallucinated text. Dialogue System facilitates human-computer interaction through natural conversation. Dialogue systems are broadly categorized into task-oriented and open-domain systems. Task-oriented systems help users complete specific goals, such as booking flights or managing appointments using structured databases [28]. Open-domain systems engage in unrestricted conversations across diverse topics [29]. Open-domain dialogue systems are prone to generating content that is not necessarily present in the input or knowledge source, leading to extrinsic hallucinations. Task-oriented systems, while more constrained, can still hallucinate due to incomplete or misclassified data [30, 31]
@@
Data to Text is a task that entails utilizing structured data, such as a table, as input to generate text that accurately and coherently represents this data as natural language text. The models of this task are prone to hallucination due to the gap between structured data and text [32]. Paraphrasing is the task of rewording text while preserving its original meaning. The model is required to generate text with semantically equivalent expressions to the source text. Hallucination can occur in this task when the model adds information not present in the source or omits crucial details. This can distort the intended message, resulting in potential misunderstandings [33]. Code Generation is an NLG task that generates code snippets based on natural language descriptions. This task is highly prone to hallucination, resulting in code that appears plausible but consists of logical, runtime, or syntax errors [34]. Such hallucinated code can lead to software malfunctions or security vulnerabilities if not identified and corrected.
@@ 4.2 Model Architecture
The second stage of the LLM development cycle involves selecting or designing an appropriate model architecture. Developers typically choose one of the transformer-based architectures for developing LLMs, such as encoder-only [44], decoder-only [6], and encoder-decoder [23]. In other scenarios, an architecture can be designed from scratch. The choice or design of an architecture is influenced by several factors, such as task complexity, available resources, and domain requirements. Once the architecture is finalized, a suitable vocabulary is constructed using tokenization methods, such as byte pair encoding (BPE) [45, 46] and SentencePiece [47]. Token embeddings are then created by mapping each vocabulary element to a numerical vector, which is initialized and fine-tuned during training.
@@ 5 Hallucination Detection
Hallucination detection involves finding instances of LLM outputs that are inaccurate, nonsensical, or inconsistent with the given input or context. Unlike traditional fact verification, which primarily verifies claims against external sources, hallucination detection involves a more comprehensive analysis. It often requires analyzing linguistic entailment and contradictions within the model’s responses. In this survey, we categorize hallucination detection methods present in the literature into retrieval, uncertainty, embedding, learning, and self-consistency-based techniques. The proposed taxonomy is illustrated in Figure 5.
@@ 5.1 Retrieval-Based Detection
Retrieval-based hallucination detection methods compare the generated output of an LLM with trusted external knowledge sources, such as databases and encyclopedias. These methods aim to ensure factual consistency and reduce the risk of unsupported content by grounding the output in verifiable information. RAG is one of the widely used retrieval-based techniques for hallucination detection [76, 77, 78]. It combines retrieval systems with generative models to fetch relevant documents and compare their content with the LLMs’ generated outputs [79]. RAG can be followed by fact-checking models to evaluate the correctness of statements [80, 79].
In order to improve retrieval-based techniques, Wang et al. [81] employed a Bayesian sequential estimation. Instead of retrieving a pre-defined number of documents, one document is retrieved at a time to assess each subclaim’s veracity. After each retrieval, the Bayesian sequential analysis decides whether to continue or stop retrieving documents. This dynamic stop-or-continue strategy optimizes the balance between retrieval costs and accurate hallucination detection, which leads to improved efficiency and precision. Building on the need for more granular analysis, Mishra et al. [82] proposed FAVA, a retrieval-augmented model trained on synthetic data to detect hallucination of various error types using span-level identification. FAVA outperformed baseline models, including GPT-4, in both span-level and binary hallucination detection, as well as factual editing tasks. KnowHalu is another improvement over retrieval-based hallucination detection [83]. It first detects non-fabrication hallucination, then performs a multi-form factual checking through step-wise reasoning, query decomposition, and knowledge retrieval. By aggregating judgments across multiple knowledge forms, KnowHalu achieves significant improvements in QA and summarization tasks by 15.7% and 5.5%, respectively.
@@ 5.2 Uncertainty-Based Detection
Probabilistic and uncertainty-based hallucination detection methods flag low-confidence outputs as potential hallucinations without requiring external knowledge retrieval. These methods exploit the inherent uncertainty in the model predictions to identify hallucinated content. The main hypothesis of uncertainty estimation is that a high uncertainty indicates that the model is guessing instead of relying on learned patterns [84, 85]. The uncertainty-based hallucination detection techniques can be classified into token-, semantic-, and supervised-based uncertainty methods. Examples show the differences between these techniques are shown in Figure 6.
@@
Token-based Approaches. Several methods have been used to quantify token-level uncertainty in LLMs to detect hallucination. Guerreiro et al. [86] detected LLM hallucination in machine translation tasks using sequence log-probability. This approach measures the model’s confidence by calculating the normalized log probability of the generated tokens. In another study [87], the log-based uncertainty measure was used to develop an uncertainty-aware framework for hallucination detection. The method relies on token probability scores derived from the LLM’s logit outputs to measure uncertainty. Instead of setting a strict threshold, uncertainty is introduced as an intermediary variable to be used adaptively. Accordingly, the model implicitly incorporates uncertainty in its decision-making. Although token-level uncertainty estimation achieved promising results, this technique is calculated based on lexical variety. Therefore, responses that give the same meaning but use different wording will not be treated as uncertain. Zhang et al. [88] enhanced uncertainty-based hallucination detection by improving token probability estimation, reducing overconfidence, and incorporating focus mechanisms. This approach focuses on the most informative keywords, unreliable tokens in historical contexts, and specific token properties.
Semantic-based Approaches. Farquhar et al. [89] proposed a semantic entropy measure to detect hallucinatory content. Semantic entropy is computed over the sentence’s meaning rather than relying on words distribution. This technique first generates multiple responses to the same prompt. These responses are then clustered based on semantic similarity. Moreover, entropy is computed across these semantic clusters. A high semantic entropy indicates that the model’s answers vary in meaning, indicating a high likelihood of hallucination, whereas low semantic entropy suggests consistent meaning across generations, which means a greater reliability. Hou et al. [90] proposed belief tree propagation (BTPROP) hallucination detection. Instead of relying solely on token-level uncertainty, BTPROP recursively decomposes a statement into logically related subclaims and builds a tree of beliefs. A hidden Markov tree is then used to integrate the model’s noisy confidence scores and logical relationships between claims. This approach robustly propagates uncertainty and corrects miscalibrated beliefs. Expanding on these directions, Chen et al. [91] introduced a graph-enhanced uncertainty modeling approach to capture the relations among entity tokens and sentences. The authors also proposed a graph-based uncertainty calibration technique that takes into account the possibility of phrase conflict with its neighbors in the semantic graph to calculate uncertainty. This method achieved a notable enhancement in passage-level hallucination detection, with an increase in the Spearman correlation by 19.78%.
@@
Supervised-based Approaches. Shelmanov et al. [92] proposed pre-trained uncertainty quantification heads to predict claim-level hallucination by leveraging attention maps and token probability features within the LLM. These transformer-based heads are trained on annotated hallucination data, which outperformed both classical unsupervised and prior supervised uncertainty methods, with strong generalization across domains and languages. Similarly, Niu et al. [93] proposed a multiple instance learning framework for hallucination detection that adaptively selects salient token embeddings that are most indicative of factual inaccuracy. The selection is guided by uncertainty metrics such as token-level and sentence-level entropy. By combining these uncertainty features with deep representation learning, this framework achieved robust and generalizable hallucination detection, outperforming prior supervised and unsupervised uncertainty-based methods.
][t] Refer to caption
@@
Figure 6: Examples of hallucination detection techniques across three uncertainty-based approaches: token-level uncertainty calculation, semantic clustering, and supervised detection via attention heads.

### 18 Mar 2026 v3

@@ 1 Introduction
Furthermore, we expanded the existing taxonomy of hallucination mitigation techniques derived from prior research [8, 9, 10] by categorizing them into four categories: prompt-, retrieval-, reasoning-, and model-centric training and adaptation-based approaches. Prompt-based mitigation approaches depend on structured prompting strategies to guide models towards generating factual content. Retrieval-based mitigation methods depend on external knowledge to ground outputs. Reasoning-based mitigation techniques, such as chain-of-thought prompting (CoT) and self-consistency, enhance logical coherence and internal consistency. Model-centric training and adaptation-based approaches involve modifying architectures, adjusting training objectives, and employing fine-tuning procedures to improve models’ intrinsic factuality and reliability. Based on our analysis, we show that no single approach completely mitigates hallucination. Consequently, more effective mitigation is a combination of complementary techniques. The most promising are hybrid approaches that combine prompting or reasoning-based techniques with retrieval-based and model-centric training and adaptation strategies.
@@
  • Hallucination mitigation discussion: The hallucination mitigation methods is classified in this survey into four main categories: prompt, retrieval, reasoning, and model-centric training and adaptation-based techniques. Each category have been discussed deeply to show its potential for hallucination mitigation.
  • Hallucination explainability discussion: The hallucination explainability methods is classified in this survey into model-derived and grounding explainability. Each category includes recent studies to show their potential for hallucination explainability.
@@ 2 Related Surveys
Given the rapid evolution of LLMs and their expanding applications across diverse domains, hallucination detection and mitigation techniques have also progressed. Consequently, there is a need to thoroughly analyze the causes of hallucinations, their detection, and their mitigation techniques to remain up to date. Investigating these factors will contribute to diagnosing the root causes of LLM hallucinations and developing more effective detection and mitigation techniques. Several surveys have been published recently on LLMs hallucination [5, 9, 8, 11, 12]. These surveys examine LLMs’ hallucination from multiple points of view and provide significant insights. Ji et al. [5] reviewed hallucination across various NLG tasks and outlined the mitigation methods applied and the evaluation metrics used in each task. Ye et al. [12] extended this work by proposing a new taxonomy for the detection and mitigation of hallucination. Zhang et al. [11] highlighted some issues related to LLMs, such as input, context, and fact-conflicting hallucinations. Tonmoy et al. [8] focused on hallucination mitigation by presenting a taxonomy that categorizes mitigation techniques into prompt engineering, retrieval-augmented generation (RAG), self-refinement, and decoding strategies. Huang et al. [9] presented a dual taxonomy of factuality and faithfulness hallucinations and defined hallucination causes in the data, training, and inference stages. The survey also linked detection and mitigation strategies to their foundational causes to guide robust system development.
@@ 3.1 Definition of Hallucination
In a psychological context, the term ”hallucination” refers to a false perception of objects or events [19]. It involves the experience of seeing, hearing, feeling, smelling, or tasting stimuli that are not actually present. These sensations are often indistinguishable from reality to the one experiencing them [20]. In the context of LLMs, hallucination refers to the generation of text that appears reasonable, fluent, and coherent but lacks grounding in factual or accurate information [21].
@@ 3.3 Hallucination in NLG Tasks
Machine Translation involves translating a text from one language to another. Models may provide grammatically accurate translations but may convey new or irrelevant meanings, particularly for low-resource languages or ambiguous inputs [24].
@@
Text Summarization is generating a concise and coherent summary that preserves the most important information from one or more source documents. In the literature, two main methods for text summarization are followed: extractive and abstractive [25]. Extractive text summarization selects and organizes the most important sentences from the source text to form a summary. Abstractive text summarization, on the other hand, paraphrases the selected sentences to form a summary. Abstractive text summarization is more prone to hallucination than extractive text summarization, since paraphrasing the text may lead to generating false information [21].
@@
Generative QA is a type of QA system that generates answers to user queries in natural language, rather than selecting a predefined answer from a database or document. It constructs its response based on the model’s comprehension of the inquiry and the provided context. Generative QA systems search for external knowledge from multiple sources to formulate a response. These sources may consist of redundant, complementary, or conflicting information [26, 27]. Therefore, it is highly susceptible to generating hallucinated text.
Dialogue System facilitates human-computer interaction through natural conversation. Dialogue systems are broadly categorized into task-oriented and open-domain systems. Task-oriented systems help users complete specific goals, such as booking flights or managing appointments using structured databases [28]. Open-domain systems engage in unrestricted conversations across diverse topics [29]. Open-domain dialogue systems are prone to generating content that is not necessarily present in the input or knowledge source, leading to extrinsic hallucinations. Task-oriented systems, while more constrained, can still hallucinate due to incomplete or misclassified data [30, 31]
@@
Data to Text is a task that entails utilizing structured data, such as a table, as input to generate text that accurately and coherently represents this data as natural language text. The models of this task are prone to hallucination due to the gap between structured data and text [32]. Paraphrasing is the task of rewording text while preserving its original meaning. The model is required to generate text with semantically equivalent expressions to the source text. Hallucination can occur in this task when the model adds information not present in the source or omits crucial details. This can distort the intended message, resulting in potential misunderstandings [33].
Code Generation is an NLG task that generates code snippets based on natural language descriptions. This task is highly prone to hallucination, resulting in code that appears plausible but consists of logical, runtime, or syntax errors [34]. Such hallucinated code can lead to software malfunctions or security vulnerabilities if not identified and corrected.
3.4 Impact of Hallucination in Critical Domains
LLMs Hallucination can pose a considerable challenge, particularly in domains where precise information is essential, such as healthcare, legal counsel, finance, and education. In such domains, LLMs hallucination can directly result into clinical harm, legal liability, and financial loss.
Hallucination in medicine is defined as the generation of an output that is not supported by authoritative clinical evidence and could alter clinical decisions [35]. One of the active research areas in employing LLMs in the medical domain is clinical documentation and consultation summarization [36]. Asgari et al. [36] found that the most common hallucination type in this domain is fabricating the planning section of clinic notes. However, detecting hallucination in this domain is challenging due to the ambiguity of the terms, which leads to diverse errors [35]. Therefore, detecting hallucination in the medical domain depends highly on the level of domain expertise and the level of detail in the prompt provided to the detection model. Domain experts are more likely to identify subtle inaccuracies in clinical terminology and reasoning, whereas non-experts may struggle to discern these errors, thereby increasing the risk of misinterpretation [35]. Accordingly, transparency, interpretability, and trustworthiness in medical AI are a necessity for Healthcare professionals to understand how and why an LLM arrived at its conclusions and thereby improve hallucination detection and mitigation in the medical domain.
Hallucination in finance raises a critical risk, since hallucinated numbers, mis-summaries of financial statements, or fabricated market events can drive misallocation of capital [37]. Various specialized LLMs, such as FinBERT [38] and BloombergGPT [39] have been developed to generate financial-related content. However, hallucination presents a major challenge to deploy such LLMs in real-world applications [37]. To mitigate hallucination in the financial domain, Kang and Liu [37] demonstrated the effectiveness of RAG with prompt-based tools in learning to generate accurate content. Moreover, Roychowdhury [40] demonstrated the effectiveness of prototyping, scaling, and LLM evolution using human feedback to minimize hallucination in the financial domain.
Hallucination in law poses a significant risk due to the authoritative tone and persuasive fluency of LLM-generated legal text, which may obscure factual or doctrinal inaccuracies. Legal hallucination refers to the generation of incorrect, fabricated, or outdated legal information, such as non-existent case law, misquoted statutes, or erroneous legal reasoning, that is not supported by valid legal sources. Such hallucinations are particularly dangerous because legal decisions often rely on precise interpretations of jurisdiction-specific laws, precedents, and procedural rules [41]. Detecting hallucination in the legal domain is particularly challenging due to the complexity and evolving nature of legal systems in different countries. Additionally, legal validity depends on temporal factors, jurisdictional scope, and contextual interpretation, all of which are difficult for LLMs to model reliably [41]. As a result, hallucination detection in legal applications requires access to up-to-date legal corpora, robust citation verification mechanisms, and domain-aware evaluation criteria that go beyond surface-level factual consistency.
Beyond model-level evaluation, hallucination detection and mitigation in critical domains must be integrated into practical system workflows. In healthcare, LLM outputs should be deployed as decision-support tools, with hallucination detection techniques serving as real-time safeguards that flag uncertain or unsupported content. Such human-in-the-loop designs align with clinical governance requirements and reduce the risk of automation bias. Similarly, in financial applications, hallucination mitigation must be coupled with auditability, confidence estimation, and traceability. This is to ensure that generated analyses and summaries can be verified against authoritative sources before influencing high-stakes decisions. Moreover, in the legal domain, LLMs must be integrated into human-in-the-loop workflows, where generated outputs serve as assistive drafts rather than final legal judgments. Human oversight by qualified legal professionals is essential to validate citations, reasoning chains, and conclusions before use in real-world settings. Moreover, regulatory and ethical considerations, including professional responsibility rules, data privacy laws, and accountability frameworks, necessitate transparency, traceability, and auditability in legal AI systems. Regulatory considerations further necessitate transparency and accountability, as hallucinated outputs may lead to legal liability or regulatory violations. Consequently, effective hallucination detection and mitigation in real-world systems requires a combination of technical safeguards, domain-aware human oversight, and compliance-aware deployment strategies.
@@ 5 Hallucination Detection
Hallucination detection involves identifying instances in LLM outputs that are inaccurate, nonsensical, or inconsistent with the input or context. Unlike traditional fact verification, which primarily verifies claims against external sources, hallucination detection involves a more comprehensive analysis. It often requires analyzing factuality and faithfulness within the model’s responses. In this survey, we categorize hallucination detection methods present in the literature into retrieval, uncertainty, embedding, learning, and self-consistency-based techniques. The proposed taxonomy is illustrated in Figure 5. Table II summarizes the hallucination detection studies. The results are mainly reported using the AUROC metric, which tests the model’s ability to distinguish between the hallucination and non-hallucination classes across all possible probability thresholds.
@@ 5.1 Retrieval-Based Detection
Retrieval-based hallucination detection methods compare an LLM’s output with trusted external knowledge sources, such as databases and encyclopedias. These methods aim to ensure factual consistency and reduce the risk of unsupported content by grounding the output in verifiable information. RAG is one of the most widely used retrieval-based techniques for hallucination detection [79, 80, 81]. It combines retrieval systems with generative models to fetch relevant documents and compare their content with LLMs-generated outputs [82]. RAG can be followed by fact-checking models to evaluate the correctness of statements [83, 82].
Ajmal et al. [84] detected hallucinations by assessing the semantic and factual consistency of generated answers against the source. The approach uses contextual similarity analysis and cross-referencing to identify discrepancies, such as incorrect entities, factual contradictions, or logically inconsistent statements. The detected discrepancies are quantified through a scoring mechanism that reflects the magnitude of hallucination. Similarly, Paudel et al. [85] introduced a hallucination detector for enterprise applications by verifying LLM outputs against the provided context, which is the retrieved document and common knowledge, which is the LLM’s parametric knowledge. It uses a modular detector (HDM-2) that checks whether generated statements contradict the provided context or conflict with common knowledge. Detection is performed at both the response level and span level, producing hallucination scores and fine-grained annotations.
In order to improve retrieval-based techniques, Wang et al. [86] used a Bayesian sequential estimation. Instead of retrieving a pre-defined number of documents, one document is retrieved at a time to assess each subclaim’s veracity. After each retrieval, the Bayesian sequential analysis decides whether to continue retrieving documents or stop. This dynamic stop-or-continue strategy optimizes the balance between retrieval costs and accurate hallucination detection, which leads to improved efficiency and precision. Building on the need for more granular analysis, Mishra et al. [87] proposed FAVA, a retrieval-augmented model trained on synthetic data to detect various types of hallucination using span-level detection. FAVA outperformed baseline models, including GPT-4, in both span-level and binary hallucination detection and in factual editing tasks. KnowHalu is another improvement over retrieval-based hallucination detection [88]. It first detects non-fabrication hallucination, then performs a multi-form factual check through step-wise reasoning, query decomposition, and knowledge retrieval. By aggregating judgments across multiple knowledge forms, KnowHalu achieved significant improvements of 15.7% and 5.5% in QA and summarization tasks, respectively. In another study, JointCQ [89] detected hallucination by extracting factual claims and generating targeted search queries from a QA pair. These queries are then used to retrieve external evidence. A verifier model checks whether each claim is supported, contradicted, or unverifiable based on the retrieved contexts. Using HalluQA, it achieved an accuracy of 80.58% and an F1-score of 83.05%.
@@ 5.2 Uncertainty-Based Detection
Probabilistic and uncertainty-based hallucination detection methods flag low-confidence outputs as potential hallucinations without requiring external knowledge retrieval. These methods exploit the inherent uncertainty in the model predictions to identify hallucinated content. The main hypothesis in uncertainty estimation is that high uncertainty indicates that the model is guessing rather than relying on learned patterns [90, 91]. Uncertainty-based hallucination detection techniques can be classified into token-, semantic-, and structure-based methods. Examples of the differences between these techniques are shown in Figure 6.
@@
Token-based Approaches. Several methods have been used to quantify token-level uncertainty in LLMs to detect hallucination. Guerreiro et al. [92] detected LLM hallucination in machine translation tasks using sequence log-probability. This approach measures the confidence of the model by calculating the normalized log probability of the generated tokens. In another study [93], the log-based uncertainty measure was used to develop an uncertainty-aware framework for hallucination detection. The method relies on token probability scores derived from LLM logit outputs to measure uncertainty. Instead of setting a strict threshold, uncertainty is introduced as an intermediary variable to be used adaptively. Accordingly, the model implicitly incorporates uncertainty in its decision-making. Although token-level uncertainty estimation achieved promising results, this technique is calculated based on lexical variety. Therefore, responses that give the same meaning but use different words will not be treated as uncertain. Zhang et al. [94] enhanced uncertainty-based hallucination detection by improving token probability estimation, reducing overconfidence, and incorporating focus mechanisms. This approach focuses on the most informative keywords, unreliable tokens in historical contexts, and specific token properties. Instead of relying on raw probability, Dasgupta et al. [95] proposed HalluShift, which detects hallucinations by measuring distribution shifts in internal hidden states and attention layers.
@@
More recent work improves upon this idea by learning to model token-level uncertainty from internal representations rather than relying only on raw probability values. Instead of treating all tokens equally, these methods train auxiliary models to identify which token-level signals, such as hidden states or probability distributions, are most indicative of hallucination. Shelmanov et al. [96] proposed pre-trained uncertainty quantification heads to predict claim-level hallucination by leveraging attention maps and token probability features within the LLM. These transformer-based heads are trained on annotated hallucination data, which outperformed both classical unsupervised and prior supervised uncertainty methods, with strong generalization across domains and languages. Similarly, Niu et al. [97] proposed a multiple instance learning framework for hallucination detection that adaptively selects salient token embeddings that are most indicative of factual inaccuracy. The selection is guided by uncertainty metrics such as token-level and sentence-level entropy. By combining these uncertainty features with deep representation learning, this framework achieved robust and generalizable hallucination detection, outperforming prior supervised and unsupervised uncertainty-based methods. HaluNet [98] detected hallucinations using internal uncertainty signals extracted in a single forward pass. It combines token log-likelihoods (confidence), entropy (distributional uncertainty), and hidden-state embeddings (semantic uncertainty) through a lightweight multi-branch neural architecture. These uncertainty signals are fused with attention-based weighting to learn how different types of uncertainty correlate with hallucination. Unlike self-consistency methods, it does not require multiple generations, making it efficient and suitable for real-time QA systems.
Semantic-based Approaches. Farquhar et al. [99] proposed a semantic entropy measure to detect hallucinatory content. Semantic entropy is computed over the sentence’s meaning rather than relying on the words’ distribution. This technique first generates multiple responses to the same prompt. These responses are then clustered based on semantic similarity. Moreover, entropy is computed across these semantic clusters. A high semantic entropy indicates that the model’s answers vary in meaning, indicating a high likelihood of hallucination, while a low semantic entropy suggests a consistent meaning across generations, which means greater reliability.
Structured-based Approaches. Recent studies proposed structured uncertainty detection, which detects hallucination by modeling the logical or semantic dependencies between claims as a graph or tree and propagating uncertainty across the connected units. Hou et al. [100] proposed detecting hallucination by belief tree propagation (BTPROP). Instead of relying solely on token-level uncertainty, BTPROP recursively decomposes a statement into logically related subclaims and builds a tree of beliefs. A hidden Markov tree is then used to integrate the model’s noisy confidence scores and logical relationships between claims. This approach robustly propagates uncertainty and corrects miscalibrated beliefs. Expanding on these directions, Chen et al. [101] introduced a graph-enhanced uncertainty modeling approach to capture the relations among entity tokens and sentences. The authors also proposed a graph-based uncertainty calibration technique that accounts for the possibility of phrase conflicts with neighbors in the semantic graph when calculating uncertainty. This method achieved a notable improvement in passage-level hallucination detection, increasing the Spearman correlation by 19.78%.
Beyond semantic graphs, recent work explores structural signals directly derived from internal attention patterns. Lookback Lens [102] is an unsupervised approach that analyzes attention for hallucination detection. By measuring the lookback ratio, which quantifies how much the model attends to prior context versus its own outputs, this approach identifies contextual hallucinations without requiring labeled data. This method achieved competitive performance with supervised detectors like RIPA and supports real-time mitigation during decoding, though it incurs higher inference time and relies on effective sampling. LapEigvals [103] modeled attention maps using a Laplacian graph and spectral features. It achieved superior results, with an AUC-ROC of 88.9% on the TriviaQA dataset. Likewise, HalluZig [104] detected hallucinations by training a supervised classifier on topological features extracted from attention dynamics. It models layer-wise attention matrices as graphs and applies zigzag persistence from topological data analysis to capture how attention structures evolve across layers. Hallucinations are identified when these signatures indicate unstable or short-lived attention patterns associated with flawed reasoning. Similarly, TOHA [105] detected hallucinations by analyzing the topological structure of attention maps inside the LLM. It builds attention graphs that links between the prompt and response tokens. Then, they compute a topological divergence score (MTop-DivG) that measures how structurally different the response is from the prompt. The method identifies a small set of hallucination-aware attention heads whose divergence consistently separates hallucinated from grounded outputs. The final detection is performed by averaging the divergence scores across these heads, which makes the method computationally efficient.
@@
Figure 6: Examples of hallucination detection techniques across three uncertainty-based approaches: token-level uncertainty calculation, semantic clustering, and structured detection via attention heads.